# WOMD-Reasoning: A Large-Scale Dataset and Benchmark for Interaction and Intention Reasoning in Driving

### v2

@@ 4 WOMD-Reasoning Dataset
4.1 Dataset Overview
Table 2 provides the statistics of Q&As in WOMD-Reasoning dataset, as well as in every Q&A class. Our Q&As can be classified into 5 categories: The first 3 categories are Factual contents: Map Environment, which includes the existence and category of intersections, existence and places of stop signs and crosswalks, and count of lanes, etc. Ego Agent’s Motion Status and Other (Surrounding) Agents’ Motion status, which describe each agent’s speed, acceleration, direction, related traffic light and sign, and relative position to the intersection center as well as to the ego agent. The other 2 categories are Reasoning contents: Interaction, which includes a summary of interactions between each surrounding agent and the ego agent and the reasoning behind such interactions. Intention, which sums up all interaction information to give the predicted intention of the ego agent, considering its responses to all the interactions. The factual description of the map environment and agents’ motion status in the dataset are good for training language scene describers, while the interaction and intention reasoning parts can be used for training language-based prediction and planning agents.
We claim that to the best of our knowledge, WOMD-Reasoning is the largest real-world language dataset for driving. As shown in Table 1, we provide roughly 3 million Q&A pairs for 63k scenes in WOMD, which are 6-113 times greater than those of other datasets we compare to. Furthermore, WOMD-Reasoning contains 346k unique traffic rule-induced and intention-induced interaction Q&As, as well as 63k comprehensive intention predictions of the ego agent containing responses to the interactions, which rarely exist in previous datasets. We will show these features through examples and statistics in the following section.
Table 2: Summary of dataset split and the quantitative distribution across different categories in the WOMD-Reasoning dataset.
@@ 4.1 Dataset Overview
4.2 Dataset Quality
@@
Figure 2: A demonstration of Q&As in each part of WOMD-Reasoning dataset and an example for human intention-induced interaction. We show Q&As in all categories regarding the scenario while demonstrating language analysis of overtaking, a human intention-induced interaction in WOMD-Reasoning dataset.
Figure 3: The Motion-LLaVA pipeline, which fine-tunes a multi-modal model with WOMD-Reasoning. Motion data go through pre-trained motion vector encoders from Multipath++ [51] and a projector layer to serve together with the questions in WOMD-Reasoning as the inputs for the LLM. The answers in WOMD-Reasoning serve as the supervision.
Table 3: Preliminary Human Evaluation of the Dataset. Note that we ask evaluators to comment on whether main part of the answer is correct, where minor errors would not be considered as wrong.
Q&A Class Map Ego Agent Other Agents Interactions Intentions
Total Q&As 38 57 339 57 10
Ave. ‘Accuracy’ 80.70% 95.32% 93.51% 94.73% 90.00%
Std. Error 6.08% 2.03% 3.08% 3.51% 10.00%
@@
To grab a sense of the accuracy of WOMD-Reasoning, we also execute a preliminary human evaluation. Consent was obtained from all the people interviewed. Note that here we ask evaluators to think a Q&A pair as correct if the correct main answer to the question is included in the answer, so the main answer is reasonably based on the human evaluators’ interpretation of the traffic data. This also does not exclude minor errors (like errors in detailed numbers or very far-away agents), since such errors are hardly distinguishable by the human evaluators. We invite 3 people to judge all Q&As from 11 consecutive scenes of the dataset, which contains 501 Q&As in total. The results are shown in Table 3. We conclude that considering our dataset is fully automatically generated, it achieves a satisfying accuracy, or at least it would significantly reduce the human labor necessary to build a more reliable dataset later. Further human verification of the dataset is still ongoing.
@@ 4.3 Vision Extension with Simulations
To support applications requiring visual information, we provide an API to extend WOMD-Reasoning scenario descriptions to the vision domain, compensating for the absence of raw camera data in the original WOMD dataset. We achieve this by utilizing an open-source traffic scenario simulation platform, namely ScenarioNet [25], and employing the MetaDrive [24] simulator to replay WOMD-Reasoning scenarios in both BEV layout and ego-view images through 3D rendering. This approach enables visual inputs to support the application of WOMD-Reasoning for future research in the vision and language domain. For video generation, we produce a 90-frame 10 Hz video aligned with Waymo’s trajectory data, including 10 frames of historical footage and 80 frames of future trajectory. Users may choose to use the historical portion for prediction tasks.
4.4 Comparison to Existing Datasets
Comparing WOMD-Reasoning with existing datasets, as shown in Table 1, we are the largest real-world dataset in terms of the total scenes, total Q&As, and the total interactions included. Also, WOMD-Reasoning supports Q&As in scene descriptions, prediction, and planning. Together with the simulated video of each covered scenario, it is suitable for training LLMs to perform nearly all major tasks in autonomous driving. Diving even deeper into the interaction reasoning in these datasets, we summarize that most interactions in previous datasets are caused by a very near distance. However, as we show in Sec. 4.2, many interactions happen when the two counterparts are initially far away, but they establish interactions by following the traffic rules or by human intentions. Our WOMD-Reasoning contains a wealth of this interaction information, as shown in Figure 4, the vocabulary statistics of the interaction and intention part of our dataset.
5 Experiments with Motion-LLaVA
@@
To demonstrate the benefits of WOMD-Reasoning in applications, in this section, we show the performances of Motion-LLaVA, a multi-modal model fine-tuned on WOMD-Reasoning. We show that the model gains significant abilities from WOMD-Reasoning in answering both factual and reasoning questions related to driving, which is supported by both qualitative showcases of its answers to driving-related questions and quantitative evaluations of language metrics.
5.1 Model Setup
We take LLaVA-v1.5-7b [28] as the pre-trained VLM. Since MultiPath++ [51] did not release their codes nor checkpoints, we pre-train the implementations by [21]. Our multi-modal fine-tuning takes 2 GPU days ($\approx$ 1 day on 2xNVIDIA A6000 GPUs) to train 1 epoch on the entire training set of WOMD-Reasoning. For evaluation, we randomly select $\approx$ 1,000 QA pairs from the validation set of WOMD-Reasoning.
5.2 Qualitative Results
@@
We present a set of Q&A examples answered by Motion-LLaVA covering each category of WOMD-Reasoning in one specific driving scenario, shown in Figure 5. We observe that Motion-LLaVA model is capable of providing accurate narratives and analysis on diverse driving-related contents thanks to the wide coverage of WOMD-Reasoning.
As WOMD-Reasoning is proposed to focus on the interaction analysis, we further present a few interaction analysis Q&As by Motion-LLaVA in Figure 6, in various interaction scenarios. These results further confirm the capacity of Motion-LLaVA for analyzing interactions.
5.3 Benchmarks of Multi-modal Reasoning Models
To comprehensively benchmark our proposed Motion-LLaVA model and fine-tuned LLaVA, we utilize a suite of standard language metrics, including BLEU-4 [35], METEOR [3], CIDEr [52], SPICE [1], and ROUGE-L [27], along with GPT Score [38] to assess the quality of the generated content. Details on the prompts used to calculate the GPT Score can be found in Appendix A.3. We present evaluation metrics of all generated answers, as well as those in the two categories, i.e. the Facts which include the road environment and agents’ motion status, and the Reasonings which cover analysis on interactions and intentions. Benchmark results are shown in Table 4.
Figure 4: Selected vocabulary statistics of WOMD-Reasoning. We stat vocabularies strongly related to traffic rule-induced and human intention-induced interactions in the interaction and intention part of WOMD-Reasoning, illustrating that our dataset contains abundant such interaction descriptions and reasoning.
@@ 5.3 Benchmarks of Multi-modal Reasoning Models
Non-Fine-tuned LLaVA [28] 0.512 0.211 0.275 1.36 0.455 2.31
Fine-tuned LLaVA [28] 0.779 0.581 0.439 5.51 0.735 6.88
@@
Only Non-Fine-tuned LLaVA [28] 0.561 0.322 0.326 1.50 0.483 2.46
Fine-tuned LLaVA [28] 0.821 0.702 0.495 6.05 0.783 6.91
@@
Only Non-Fine-tuned LLaVA [28] 0.262 0.097 0.201 0.06 0.283 0.91
Fine-tuned LLaVA [28] 0.596 0.452 0.357 2.26 0.557 6.14
@@
We first verify the effectiveness of these language metrics with shuffled ground-truths, which attach shuffled true answers to questions randomly. Results confirm that low scores are granted for this model, indicating that a similar structure with wrong answer content would not be able to cheat our metrics. Then we benchmark the off-the-shelf LLaVA [28], where BEV plots are used as the vision input for the driving scenario. This non-fine-tuned version earns a very low score in all metrics, confirming that LLaVA itself cannot answer driving-related questions correctly.
Then, we perform fine-tuning on WOMD-Reasoning. We compare two models with different modalities as motion input: LLaVA using the BEV plot of the driving scene as the vision input [28], and Motion-LLaVA which uses motion data as input. We observe that both fine-tuned multimodal models significantly outperform previous baselines across all metrics and question types, underscoring the value of fine-tuning on WOMD-Reasoning for addressing driving-related queries. Compared to the fine-tuned LLaVA, Motion-LLaVA achieves superior performance across all metrics and question types, highlighting the advantages of utilizing the motion data as input. We hypothesize that this performance gap is due to the information loss that occurs when converting motion data into a BEV plot with a scale bar. We also visualize and compare the answers of Motion-LLaVA and these baselines on the same driving scene in Appendix A.4.
Figure 6: An example set of answers to interaction-related questions generated by Motion-LLaVA on various WOMD scenarios. We see that Motion-LLaVA is capable to perform diverse interaction reasonings.
In addition to the language metrics, we provide other objective metrics to evaluate answers to quantitative questions regarding Facts. Of the 1,000 questions in the factual test set, 130 are quantitatively assessable (e.g., position, speed, etc.). For these, we report the root median squared error (Median Error), and accuracy within a 1.0-meter tolerance (Acc1.0), following similar metrics used in [19] as shown in Table 6. Results confirm that Motion-LLaVA is capable of providing accurate answers to factual questions.
None 0.814 0.684 0.491 5.95 0.774 30.0% 2.00 6.79
Motion Caption 0.824 0.711 0.498 6.05 0.775 42.6% 2.24 6.97
AutoEncoder 0.837 0.727 0.516 6.35 0.803 47.3% 1.41 7.07
Motion Prediction (Ours) 0.840 0.736 0.516 6.35 0.794 56.2% 1.00 7.09
Table 6: Accuracy of answers to factual questions.
@@
Fine-tuned LLaVA [28] 10.9% 5.0
@@
5.4 Ablation on the Design of Motion-LLaVA
Chain-of-Thought Inference. To address the hallucinated response problem [15, 49, 45] in reasoning interactions, we use a Chain-of-Thought (CoT) approach to provide Motion-LLaVA with structured context for more accurate reasoning. Table 7 compares model performance with and without CoT on interaction and intention analysis. We see that with CoT, the model performs better in GPT Scores. We further show that if our motion data translation program A.1.1 is used during inference to provide contexts, i.e., with factual description context (an example can be found in Appendix A.6) and CoT, the performance can be significantly improved, proving the potential of the CoT inference strategy.
Influence of Pre-training for the Motion Vector Encoder. To establish a pre-training strategy for the motion vector encoder, we examine factual question answering across Motion-LLaVA with four pre-training configurations: no pre-training, autoencoder-style pre-training, motion-caption pre-training [7], and pre-training with a motion prediction task. All configurations shared the same encoder structure. In the motion-caption pre-training, we pre-train all but LLM weights in Motion-LLaVA, by penalizing errors in captioning the motion. For autoencoder-style pre-training and pre-training with a motion prediction task, we retain only the encoder after pre-training, to function as the motion vector encoder for Motion-LLaVA.
@@ 6 Conclusion
We first build WOMD-Reasoning, a language annotation dataset upon WOMD, and then we propose Motion-LLaVA, an LLM fine-tuned on it, both focusing on analyzing and reasoning interactions. Specifically, our WOMD-Reasoning emphasizes interactions induced by traffic rules and human intentions, such as interactions governed by traffic lights and signs, and interactions like overtaking and following, which are both rarely covered in previous datasets. Besides, WOMD-Reasoning provides 3 million Q&A pairs, which makes it the largest real-world language dataset in autonomous driving to the best of our knowledge, covering a range of topics from scene descriptions, prediction, to planning. Boosted by WOMD-Reasoning, we benchmark a set of methods, to propose Motion-LLaVA, which fine-tunes a multi-modal model to accurately answer questions covering both driving scenario descriptions and interaction reasonings. Motion-LLaVA proves the effectiveness of WOMD-Reasoning in assisting real-world driving through various language-involved downstream tasks.
Table 7: Effects of CoT in answering reasoning-related questions.
Direct 0.616 0.474 0.365 2.54 0.578 6.59
CoT 0.614 0.474 0.366 2.52 0.571 6.76
CoT w/ Facts Contexts 0.664 0.542 0.403 3.10 0.631 7.64
Discussions. Although WOMD-Reasoning demonstrates its effectiveness in case studies, human evaluations, and multi-modal fine-tuning, we only manually verified a small subset. With human verification ongoing, the dataset will achieve even greater reliability. Also, WOMD-Reasoning is subject to the limitations of WOMD, e.g. its accuracy and lack of vision modal. We leverage a simulation approach to address the lack of support for video-Q&A, while acknowledging that it still exhibits a domain gap from real-world visual input. As WOMD continues to release additional modalities such as camera features and lidar data recently, and as the iterative development of the simulator [25, 24] we use, these limitations are expected to be naturally addressed.

### v3

@@ 4 WOMD-Reasoning Dataset Specifications
4.1 Dataset Statistics
Table 2 provides the statistics of Q&As in WOMD-Reasoning dataset, as well as in every Q&A class. Our Q&As can be classified into two parts: The Factual contents and the Reasoning contents, where the first part serves as the context for the second part. The factual contents contain three topics: Map Environment, which includes the existence and category of intersections, existence and places of stop signs and crosswalks, and count of lanes, etc. Ego Agent’s Motion Status and Other (Surrounding) Agents’ Motion status, which describe each involved agent’s speed, acceleration, direction, related traffic light and sign, and relative position to the intersection center as well as to the ego agent. The factual descriptions alone are also good for training language scene describers.
Our key part, the Reasoning contents, first comes with Q&As on Interactions, which includes a summary of interactions between each surrounding agent and the ego agent and the reasoning behind such interactions. Then by summing up all interaction information, we offer the Intentions, which give the predicted intention of the ego agent, considering its responses to all the interactions. The interaction and intention reasoning parts provide an unprecedented tool for training language-based prediction and planning agents.
As shown in Table 1, WOMD-Reasoning contains 346k unique traffic rule-induced and intention-induced interaction Q&As, as well as 63k comprehensive intention predictions of the ego agent containing responses to the interactions, which rarely exist in previous datasets. We will show these features through examples and statistics in the following section. Furthermore, we also claim that to the best of our knowledge, WOMD-Reasoning is the largest real world language data set for driving, since our WOMD-Reasoning provides roughly 3 million Q&A pairs for 63k scenes in WOMD, which are 6 - 113 times greater than those of other datasets we compare to.
Compared with existing data sets, as shown in Table 1, our WOMD-Reasoning is the largest real-world dataset in terms of total scenes, total Q&As, and the total interactions included. Also, WOMD-Reasoning supports Q&As in scene descriptions, prediction, and planning. Together with the simulated video of each covered scenario, it is suitable for training VLMs to perform nearly all major tasks in autonomous driving. Diving even deeper into the interaction reasoning in these datasets, we summarize that most interactions covered in previous datasets are caused by a very near distance. However, as we show in Sec. 4.3, many interactions happen when the two counterparts are initially far away, but they establish interactions by following the traffic rules or by human intentions. Our WOMD-Reasoning contains a wealth of this interaction information, as shown in Figure 2, the vocabulary statistics of the interaction and intention part of our dataset.
Table 2: Sizes of WOMD-Reasoning dataset.
@@ 4.1 Dataset Statistics
Figure 2: Selected vocabulary statistics in WOMD-Reasoning. We stat vocabularies strongly related to traffic rule-induced and human intention-induced interactions in WOMD-Reasoning, illustrating that it contains abundant such interaction descriptions and reasoning.
Table 3: Human Evaluations of WOMD-Reasoning Dataset.
Q&A Class All Interactions Intentions
Accuracy 91.99% 91.03% 87.50%
4.2 Interaction Contents
@@
Figure 3: A demonstration of Q&As in each part of WOMD-Reasoning dataset. We show Q&As in all categories regarding the scenario while demonstrating language analysis of overtaking, a human intention-induced interaction in WOMD-Reasoning dataset.
Figure 4: The Motion-LLaVA pipeline fine-tuning a multi-modal model with WOMD-Reasoning. Motion data go through pre-trained motion vector encoders from Multipath++ (Varadarajan et al., 2021) and a projector layer to serve together with the questions in WOMD-Reasoning as the inputs. The answers in WOMD-Reasoning serve as the supervision.
@@
4.3 Dataset Quality Evaluation by Human
To grab a sense of the accuracy of WOMD-Reasoning, we execute a human evaluation process. Consent was obtained from all the people interviewed. We ask interviewers to evaluate whether the correct answer to the question is included in the answer provided in the dataset. We invite 4 people to judge 1,610 Q&As in total. The results are shown in Table 3. We conclude that WOMD-Reasoning achieves a satisfying accuracy, especially considering it is automatically generated. Further human-based data cleaning is ongoing.
@@ 4.4 Vision Extension with Simulations
To support applications requiring visual information, we provide an API to extend WOMD-Reasoning scenario descriptions to the vision domain, compensating for the absence of raw camera data in the original WOMD dataset. We achieve this by utilizing an open-source traffic scenario simulation platform, namely ScenarioNet (Li et al., 2023), and employing the MetaDrive (Li et al., 2022) simulator to replay WOMD-Reasoning scenarios in both BEV layout and ego-view images through 3D rendering. This approach enables visual inputs to support the application of WOMD-Reasoning for future research in the vision and language domain. For video generation, we produce a 90-frame 10 Hz video aligned with Waymo’s trajectory data, including 10 frames of historical footage and 80 frames of future trajectory. Users may choose to use the historical portion for prediction tasks.
5 Dataset Application and Evaluation with Motion-LLaVA
@@
To demonstrate a real-world application of WOMD-Reasoning, we design Motion-LLaVA, a multi-modal model taking motion data as input, and fine-tune it on WOMD-Reasoning. We show that the model gains significant abilities from WOMD-Reasoning in predicting interactions, especially those caused by traffic rules, and in planning right-of-way compliant route for the ego. Furthermore, its abilities extend to answer various questions related to driving across factual and reasoning categories, which is supported by both qualitative showcases and quantitative evaluations on language metrics. These highlight the broad applications of WOMD-Reasoning for the autonomous driving industry.
5.1 Interaction Prediction
Since WOMD-Reasoning is proposed to focus on the interaction analysis, one of the key abilities of Motion-LLaVA fine-tuned on it is to predict interactions, especially the traffic rule-induced interactions. We present a few interaction prediction examples answered by Motion-LLaVA in Figure 5. These results confirm that by fine-tuning on WOMD-Reasoning, Motion-LLaVA unleashes the potential of language models for predicting traffic rule-related interactions in unseen real-world scenarios.
Figure 5: Interaction predictions made by Motion-LLaVA on various WOMD scenarios.
5.2 Traffic Rule Compliant Planning
Based on thorough interaction analysis, traffic rule compliant planning can naturally be achieved by summarizing interactions, as we do in producing ego agents’ intentions in WOMD-Reasoning. Motion-LLaVA also perfectly learns this ability. In Figure 6, we demonstrate that Motion-LLaVA fine-tuned on WOMD-Reasoning can offer traffic rule-compliant future trajectory plans, which clearly extends the applications of WOMD-Reasoning into planning.
Figure 6: Traffic rule-compliant planning made by Motion-LLaVA on various WOMD scenarios.
5.3 Answering Various Driving-related Questions
@@
We further present a set of Q&A examples answered by Motion-LLaVA covering each category of WOMD-Reasoning in one specific driving scenario, shown in Figure 7. We observe that Motion-LLaVA model is capable of providing accurate narratives and analysis on diverse driving-related contents thanks to the wide coverage of WOMD-Reasoning.
5.4 Validation of Motion-LLaVA
To validate the effectiveness of Motion-LLaVA and the fine-tuning process, we present evaluation metrics of all generated answers, as well as those in the two categories, i.e. the Facts which include the road environment and agents’ motion status, and the Reasonings which cover analysis on interactions and intentions. Benchmark results are shown in Table 4. We first verify the effectiveness of these language metrics with shuffled ground-truths, which attach shuffled true answers to questions randomly. Results confirm that low scores are granted for this model, indicating that a similar structure with wrong answer content would not be able to cheat our metrics. Then we benchmark the off-the-shelf LLaVA (Liu et al., 2024), where BEV plots are used as the vision input for the driving scenario. This non-fine-tuned version earns a very low score in all metrics, confirming that LLaVA itself cannot answer driving-related questions correctly. Then, we perform fine-tuning on WOMD-Reasoning. We compare two models with different modalities as motion input: LLaVA using the BEV plot of the driving scene as the vision input (Liu et al., 2024), and Motion-LLaVA which uses motion data as input. We observe that both fine-tuned multi-modal models significantly outperform previous baselines across all metrics and question types, underscoring the value of fine-tuning on WOMD-Reasoning for addressing driving-related queries. Compared to the fine-tuned LLaVA, Motion-LLaVA achieves superior performance across all metrics and question types, highlighting the advantages of utilizing the motion data as input. We hypothesize that this performance gap is due to the information loss that occurs when converting motion data into a BEV plot with a scale bar. We also visualize and compare the answers of Motion-LLaVA and these baselines on the same driving scene in Appendix A.6.
@@ 5.4 Validation of Motion-LLaVA
Non-Fine-tuned LLaVA (Liu et al., 2024) 0.512 0.211 0.275 1.36 0.455 2.31
Fine-tuned LLaVA (Liu et al., 2024) 0.779 0.581 0.439 5.51 0.735 6.88
@@
Only Non-Fine-tuned LLaVA (Liu et al., 2024) 0.561 0.322 0.326 1.50 0.483 2.46
Fine-tuned LLaVA (Liu et al., 2024) 0.821 0.702 0.495 6.05 0.783 6.91
@@
Only Non-Fine-tuned LLaVA (Liu et al., 2024) 0.262 0.097 0.201 0.06 0.283 0.91
Fine-tuned LLaVA (Liu et al., 2024) 0.596 0.452 0.357 2.26 0.557 6.14
@@
In addition to language metrics, we also provide objective metrics to the answers to questions regarding Facts. We report the root median squared error (Median Error), and accuracy within a 1.0-meter tolerance (Acc1.0) in Table 5, following similar metrics used in (Jin et al., 2023). Results confirm that Motion-LLaVA is capable of providing numerically accurate answers to factual questions.
Table 5: Accuracy of Motion-LLaVA answers to factual questions.
@@
Fine-tuned LLaVA (Liu et al., 2024) 10.9% 5.0
@@
6 Dataset Effectiveness in Vehicle Trajectory Prediction
In this section, we evaluate the effectiveness of WOMD-Reasoning in a real-world driving task - vehicle trajectory prediction. Our experiments indicate that with our WOMD-Reasoning dataset involved, the metrics of vehicle trajectory predictions are significantly improved.
6.1 Language-assisted Trajectory Prediction Experiment Setup
We adapt the Multipath++ (Varadarajan et al., 2021) trajectory prediction model to accommodate the language input. The adaption mainly contains two steps:
Language Acquisition and Encoding The language we use comes from Motion-LLaVA’s output. We run Motion-LLaVA on the cases covered by WOMD-Reasoning dataset in the training and validation set of WOMD. For each scenario included, Motion-LLaVA would provide a comprehensive set of Q&As, where we pick the interaction Q&As for this experiment. Each of these Q&As includes the interaction info between a specific agent and the ego agent. We pick the answer parts of these Q&As to use. These answers are then fed into a T5 encoder (Raffel et al., 2023) followed by a few MLP layers to be encoded. In the end, for each scenario included, we have a set of language embeddings, each contains info on the interaction between the ego agent and one specific agent.
Introducing Languages into Multipath++ Our next step is to introduce these encoded languages into each agent’s feature. For Multipath++, we use MPA (Konev, 2022), an open-source Multipath++ implementation, as our code base. Our language introduction module is inserted right after their "Other Agents History Encoder", and before their "MCG (multi-context gating) encoder". Our module is a cross-attention block, which takes each agent’s history encoding as queries, letting them cross-attend to their corresponding language embeddings containing info about the interaction between the ego agent and that specific agent. In this way, the interaction information in Motion-LLaVA is introduced into the features of each agent.
6.2 Enhancing Trajectory Prediction with Motion-LLaVA Outputs
Table 6: Benefits of Motion-LLaVA Outputs for Prediction Task.
Model $minFDE_{6}$ $MR_{6}$
Multipath++ (Konev, 2022) 1.27 12.59
+ Motion-LLaVA Outputs 1.18 11.69
$\Delta$7.35% -7.10%
As shown in Table 6, we observe significant metric improvements with language involved, which strongly supports WOMD-Reasoning’s ability to help downstream tasks like predictions. It can also help to understand the driving behaviors, and to improve explainability of vehicle trajectory prediction models.
@@ 7 Conclusion
We first build WOMD-Reasoning, a language-centered multi-modal dataset upon WOMD, focusing on analyzing and reasoning interactions. Specifically, our WOMD-Reasoning emphasizes interactions induced by traffic rules and human intentions, such as interactions governed by traffic lights and signs, and interactions like overtaking and following, which are both rarely covered in previous datasets. Besides, WOMD-Reasoning provides 3 million Q&A pairs, which makes it the largest real-world language dataset in autonomous driving to the best of our knowledge, covering a range of topics from scene descriptions, prediction, to planning. To demonstrate the applications of WOMD-Reasoning, we propose Motion-LLaVA, which fine-tunes a multi-modal model to accurately answer questions covering both driving scenario descriptions and interaction reasonings. Motion-LLaVA realizes the function of interaction prediction and traffic rule-compliant planning, taking advantages of the WOMD-Reasoning and the text-based knowledge intrinsically built in language models. Motion-LLaVA proves the effectiveness of WOMD-Reasoning in assisting real-world driving through supporting various language-involved downstream tasks.